# The Unbounded Denominators Conjecture

### v3

@@ Abstract.
We prove the unbounded denominators conjecture in the theory of noncongruence modular forms for finite index subgroups of $\mathrm{SL}_{2}(\mathbf{Z})$. Our result includes also Mason’s generalization of the original conjecture to the setting of vector-valued modular forms, thereby supplying a new path to the congruence property in rational conformal field theory. The proof involves a new arithmetic holonomicity bound of a potential-theoretic flavor, together with Nevanlinna’s second main theorem, the congruence subgroup property of $\mathrm{SL}_{2}(\mathbf{Z}[1/p])$, and a close description of the Fuchsian uniformization $D(0,1)/\Gamma_{N}$ of the Riemann surface $\mathbf{C}\smallsetminus\mu_{N}$.
F.C. was supported in part by NSF Grant DMS-2001097.
@@ 1. Introduction
Let $N$ be any positive integer, and let $f(\tau)\in\mathbf{Z}\llbracket q^{1/N}\rrbracket$ for $q=\exp(\pi i\tau)$ be a holomorphic function on the upper half plane. Suppose there exists an integer $k$ and a finite index subgroup $\Gamma\subset\mathrm{SL}_{2}(\mathbf{Z})$ such that
$f\left(\frac{a\tau+b}{c\tau+d}\right)=(c\tau+d)^{k}f(\tau),\quad\quad\forall\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\in\Gamma,$
@@
possessing a formal Puiseux branch in $\mathbf{Z}\llbracket\lambda(\tau/m)/16\rrbracket\otimes\mathbf{C}$ for some $m\in\mathbf{N}$ are exactly the congruence coverings $Y_{\Gamma}=\mathbf{H}/\Gamma\to\mathbf{H}/\Gamma(2)=Y(2)$, with $\Gamma$ ranging over all congruence subgroups of $\Gamma(2)$. The reverse implication follows from [Shi71, Theorem 3.52], and reflects the fact that the $q$-expansions of eigenforms on congruence subgroups are determined by their Hecke eigenvalues (see also [Kat73, §1.2]).
@@
The resulting Fourier coefficients are closely linked to the dimensions of the irreducible representations of the exceptional Lie group $E_{8}(\mathbf{C})$, and in particular they are integers. To be more precise: the modular function $j^{1/3}$ coincides with the graded dimension of the level one highest-weight representation of the affine Kac–Moody algebra $E_{8}^{(1)}$; see Gannon’s book [Gan06, page 196] for a broad view on this topic and its relation to mathematical physics. The unbounded denominators conjecture (Theorem 1.0.1) now implies that $j^{1/3}$ must be a modular function on a congruence subgroup, and indeed one may easily confirm that it is a Hauptmodul for the level $3$ subgroup which is the kernel of the composite $\mathrm{PSL}_{2}(\mathbf{Z})\rightarrow\mathrm{PSL}_{2}(\mathbf{F}_{3})=A_{4}\rightarrow\mathbf{Z}/3\mathbf{Z}$. One final example is the function
$h:=\frac{\lambda(\tau)(1-\lambda(\tau))}{16}=\left(\frac{\eta(\tau/2)\eta(2\tau)}{\eta(\tau)^{2}}\right)^{24}$
of level $\Gamma^{0}(2)\supset\Gamma(2)$; here the complete list of $n$ for which $h^{1/n}$ is either congruence modular or has bounded denominators are the divisors of $24$. One may even compute that the abelianization of the congruence completion $\widehat{\Gamma^{0}(2)}$ of $\Gamma^{0}(2)$ (where $\Gamma^{0}(2)$ is considered as a subgroup of $\mathrm{PSL}_{2}(\mathbf{Z})$) is $\mathbf{Z}/24\mathbf{Z}\oplus\mathbf{Z}/2\mathbf{Z}$. (The other $\mathbf{Z}/2\mathbf{Z}$ extension corresponds to $\sqrt{1-64h}=1-2\lambda$.)
In a similar vein pertaining to the examples from the representation theory of vertex operator algebras, we prove in our closing section 7 the natural generalization of Theorem 1.0.1 to components of vector-valued modular forms for $\mathrm{SL}_{2}(\mathbf{Z})$, in particular resolving — in a sharper form, in fact — Mason’s unbounded denominators conjecture [Mas12, KM08] on generalized modular forms.
@@ 1.1.1. The Diophantine principle
Lemma 1.1.1 follows upon combining the following two obsevations, fixing some $1>\eta>R^{-1}$:
@@
We shall refer to the first inequality as a Liouville lower bound, following its use by Liouville in his proof of the lower bound $|\alpha-p/q|\gg 1/q^{n}$ for algebraic numbers $\alpha\neq p/q$ of degree $n\geq 1$. We shall refer to the second inequality as a Cauchy upper bound, following the example above where it comes from an application of the Cauchy integral formula. The first non-trivial generalization of Lemma 1.1.1 was Émile Borel’s theorem [Bor94]. Dwork famously used a $p$-adic generalization of Borel’s theorem in his $p$-adic analytic proof of the rationality of the zeta function of an algebraic variety over a finite field (see Dwork’s account in the book [DGS94, Chapter 2]). The simplest non-trivial statement of Borel’s theorem is that an integral formal power series $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ must already be a rational function as soon as it has a meromorphic representation as a quotient of two convergent complex-coefficients power series on some disc $D(0,R)$ of a radius $R>1$. The subject of arithmetic algebraization blossomed at the hands of many authors, including most prominently Carlson, Pólya, Robinson, Salem, Cantor, D. &\& G. Chudnovsky, Bertrandias, Zaharjuta, André, Bost, Chambert-Loir [CL02, BCL09, Ami75], [And04, § I.5], [And89, § VIII]. A simple milestone that we further develop in our § 2 is André’s algebraicity criterion [And04, Théorème 5.4.3], stating in a particular case that an integral formal power series $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ is algebraic as soon as the two formal functions $x$ and $f$ admit a simultaneous analytic uniformization — that means an analytic map $\varphi:(D(0,1),0)\to(\mathbf{C},0)$ such that the composed formal function germ $f(\varphi(z))\in\mathbf{C}\llbracket z\rrbracket$ is also analytic on the full disc $D(0,1)$, and such that $\varphi$ is sufficiently large in terms of conformal size, namely: $|\varphi^{\prime}(0)|>1$. For example, for any integer $m$, the algebraic power series $f=(1-m^{2}x)^{1/m}\in\mathbf{Z}\llbracket x\rrbracket$ admits the simultaneous analytic uniformization $x=(1-e^{Mz})m^{-2}$ and $f=e^{Mz/m}$, where the conformal size $|\varphi^{\prime}(0)|=M/m^{2}$ can clearly be made arbitrarily large by making a suitable choice of $M$.
@@ 1.1.2. Modularity and simultaneous uniformizations of $f$ and $\lambda$
Let us now explain the relevance of arithmetic holonomy rank bounds to the unbounded denominator conjecture. After reducing to weight $k=0$ as above, the functions $f=f(\tau)$ and $x:=\lambda(\tau)/16\in q+q^{2}\mathbf{Z}\llbracket q\rrbracket$ are algebraically dependent and share both (we assume) the property of integral Fourier coefficients at the cusp $q=0$. Let us assume for the purpose of this sketch that $f(\tau)\in\mathbf{Z}\llbracket q\rrbracket$ with $q=e^{\pi i\tau}$, i.e. that the cusp $i\infty$ has width dividing $2$. Then the formal inverse series expansion
$q=x+8x^{2}+91x^{3}+\cdots\in x+x^{2}\mathbf{Z}\llbracket x\rrbracket$
of (1.0.2) has integer coefficients, expressing the identity $\mathbf{Z}\llbracket q\rrbracket=\mathbf{Z}\llbracket x\rrbracket$ of formal power series rings, and that formal substitution turns our integral Fourier coefficients hypothesis into an algebraic power series with integer coefficients: henceforth in this introductory sketch we switch to writing, by a mild and harmless notational abuse, simply $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ in place of $f(\tau)$ and $\lambda(q)$ in place of (1.0.2). In the general case of arbitrary cusp width, which we need anyhow for the inner workings of our proof even if one is ultimately interested in the $\mathbf{Z}\llbracket q\rrbracket$ case, we will only have $f\in\mathbf{Z}[1/N]\llbracket x\rrbracket$ when we write out $f(\tau)$ as a power series in $x:=\sqrt[N]{\lambda/16}$ to accommodate the Puiseux series — but there is still a hidden integrality property which we can exploit. That leads to some mild technical nuance with the power series (2.0.2) — think of $t=q^{1/N}$, $x(t)=\sqrt[N]{\lambda(t^{N})/16}$ and $p(x)=x^{N}$ — in our refinement (2.0.3) of André’s theorem.
@@ 1.1.3. A finite local monodromy leads to an overconvergence
It turns out, and this is the key to our method and already answers André’s question in [And04, Appendix, A.5], that a different choice of $\varphi(z)$ allows one to arithmetically distinguish between these two cases (algebraic and transcendental), and to have the algebraicity of $f(x)$ recognized by André’s Diophantine criterion by way of an “overconvergence.” The common feature of these two functions $f(x)$ and $F(x)$ — coming respectively out of modular forms of weights 0 and $1$ — is that they both vary holonomically in $x\in\mathbf{C}\smallsetminus\{0,1/16\}$: they satisfy linear ODEs with coefficients in $\mathbf{Q}[x]$ and no singularities [^1] With nontrivial local monodromy. The precise definition is in 2.0.4. apart from the three punctures $x=0,1/16,\infty$ of $Y(2)=\mathbf{H}/\Gamma(2)$. The difference feature is that their respective local monodromies around $x=0$ are finite for the case of $f(x)$ (a quotient of $\mathbf{Z}/N$, with the order $N$ equal to the LCM of cusp widths, or Wohlfahrt level [Woh64] of $f(x)$); and infinite for the case of $F(x)$ (isomorphic to $\mathbf{Z}$, corresponding more particularly to the fact that this particular hypergeometric function acquires a $\log{x}$ term after an analytic continuation around a small circle enclosing $x=1/16$). If now we perform the variable change $x\mapsto x^{N}$, redefaulting to $x:=\sqrt[N]{\lambda(q^{N})/16}$, that resolves the $N$-th root ambiguity in the formal Puiseux branches of $f(x)$ at $x=0$, and the resulting algebraic power series $f(x^{N})\in\mathbf{Z}\llbracket x^{N}\rrbracket\subset\mathbf{Z}\llbracket x\rrbracket$ has turned holonomic on $\mathbf{P}^{1}\smallsetminus\{16^{-1/N}\mu_{N},\infty\}$: singularities only at $16^{-1/N}\mu_{N}\cup\{\infty\}$ (but not at $x=0$: this key step of exploiting arithmetic algebraization is the same as in Ihara’s arithmetic connectedness theorem [Iha94, Theorem 1], which together with Bost’s extension [Bos99] to arithmetic Lefschetz theorems have in equal measure been inspirational for our whole approach to the unbounded denominators conjecture). Since $\lambda:D(0,1)\to\mathbf{C}\smallsetminus\{1\}$ has fiber $\lambda^{-1}(0)=\{0\}$, the function $\varphi(z):=\sqrt[N]{\lambda(z^{N})/16}:D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ is still holomorphic on the unit disc $|z|<1$, and under this tautological choice, both functions $f(x^{N})$ and $F(x^{N})$ continue to be at the borderline of André’s algebraicity criterion: $|\varphi^{\prime}(0)|=1$.
@@
But if instead of the tautological simultaneous uniformization we take $\varphi:D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ to be the universal covering map (pointed at $\varphi(0)=0$), then either by a direct computation with monodromy, or by Cauchy’s analyticity theorem on the solutions of linear ODEs with analytic coefficients and no singularities in a disc, we have both function germs $x:=\varphi(z)$ and $f(x):=f(\varphi(z))$ holomorphic, hence convergent, on the full unit disc $D(0,1)$. In contrast, now $F(\varphi(z))$ converges only up to the “first” nonzero fiber point in $\varphi^{-1}(0)\smallsetminus\{0\}$, giving a certain radius rather smaller than $1$. We must have the strict lower bound $|\varphi^{\prime}(0)|>1$, because the preceding unit-radius holomorphic map $\sqrt[N]{\lambda(z^{N})/16}:D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ has to factorize properly through the universal covering map. Indeed in Theorem 5.1.4, using an explicit description by hypergeometric functions of the multivalued inverse of the universal covering map of $\mathbf{C}\smallsetminus\mu_{N}$ based on Poincaré’s ODE approach [Hem88] to the uniformization of Riemann surfaces, we find an exact formula for this uniformization radius in terms of the Euler Gamma function. [^2] André pointed out to us that this explicit formula has previously been obtained by Kraus and Roth, see [KR16, Remark 5.1]. See also [Gol69, §III.1]. Hence the algebraicity of $f(x)$ gets witnessed by André’s criterion; and the formal new result that we get already at this opening stage (see Theorem 7.2.1) is that any integral formal power series solution $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ to a linear ODE $L(f)=0$ without singularities on $\mathbf{P}^{1}\smallsetminus\{0,1/16,\infty\}$ is in fact algebraic as soon as the linear differential operator $L$ has a finite local monodromy $\mathbf{Z}/N$ around the singular point $x=0$. More than this: the quantitative Corollary 2.0.5 proves that the totality of such $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ at a given $N$ span a finite-dimensional $\mathbf{Q}(x)$-vector space, and gives an upper bound on its dimension as a function of the Wohlfahrt level parameter $N$. Now since a (noncongruence) counterexample $f(\tau)\in\mathbf{Z}\llbracket q\rrbracket$ to Theorem 1.0.1 would not exist on its own but spawn a whole sequence $f(p\tau)\in\mathbf{Z}\llbracket q\rrbracket$ of $\mathbf{Q}(x)$-linearly independent counterexamples at growing Wohlfahrt level $N\mapsto Np$, our idea is to measure up the supply of these putative (fictional) counterexamples alongside the congruence supply at a gradually increasing level until together they break the quantitative bound (2.0.3) supplied by our arithmetic holonomy Theorem 2.0.1.

### v4

@@ Abstract.
We prove the unbounded denominators conjecture in the theory of noncongruence modular forms for finite index subgroups of $\mathrm{SL}_{2}(\mathbf{Z})$. Our result includes also Mason’s generalization of the original conjecture to the setting of vector-valued modular forms, thereby supplying a new path to the congruence property in rational conformal field theory. The proof involves a new arithmetic holonomicity bound of a potential-theoretic flavor, together with Nevanlinna second main theorem, the congruence subgroup property of $\mathrm{SL}_{2}(\mathbf{Z}[1/p])$, and a close description of the Fuchsian uniformization $D(0,1)/\Gamma_{N}$ of the Riemann surface $\mathbf{C}\smallsetminus\mu_{N}$.
F.C. was supported in part by NSF Grant DMS-2001097. Y.T. was supported in part by NSF grant DMS-2231958 and a Sloan Research Fellowship. Some of the work was done when Y.T. was at CNRS and Université Paris-Saclay from February 2020 to June 2021 and at Princeton University from July 2021 to June 2022.
@@ 1. Introduction
Let $N$ be any positive integer, and let $f(\tau)\in\mathbf{Z}\llbracket q^{1/N}\rrbracket$ for $q=\exp(\pi i\tau)$ be a holomorphic function on the upper half plane $\mathbf{H}$. Suppose there exists an integer $k$ and a finite index subgroup $\Gamma\subset\mathrm{SL}_{2}(\mathbf{Z})$ such that
$f\left(\frac{a\tau+b}{c\tau+d}\right)=(c\tau+d)^{k}f(\tau),\quad\quad\textrm{for all }\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\in\Gamma,$
@@
possessing a formal Puiseux branch in $\mathbf{Z}\llbracket\lambda(\tau/m)/16\rrbracket\otimes\mathbf{C}$ for some $m\in\mathbf{N}_{>0}$ are exactly the congruence coverings $Y_{\Gamma}=\mathbf{H}/\Gamma\to\mathbf{H}/\Gamma(2)=Y(2)$, with $\Gamma$ ranging over all congruence subgroups of $\Gamma(2)$. The reverse implication is a theorem of Shimura [Shi59] (presented in his book as [Shi71, Theorem 3.52]), and reflects the fact that the $q$-expansions of eigenforms on congruence subgroups are determined by their Hecke eigenvalues (see also [Kat73, § 1.2]).
@@
The resulting Fourier coefficients are closely linked to the dimensions of the irreducible representations of the exceptional Lie group $E_{8}(\mathbf{C})$, and in particular they are integers. To be more precise: the modular function $j^{1/3}$ coincides with the graded dimension of the level one highest-weight representation of the affine Kac–Moody algebra $E_{8}^{(1)}$; see Gannon’s book [Gan06, § 0.5 and § 3.2.3] for a broad view on this topic and its relation to mathematical physics. The unbounded denominators conjecture (Theorem 1.0.1) now implies that $j^{1/3}$ must be a modular function on a congruence subgroup. (Strictly speaking, since $j^{1/3}$ is a Laurent series rather than a power series, one applies Theorem 1.0.1 to $j^{1/3}\cdot\Delta$ and then divides through by $\Delta$.) One readily confirms that $j^{1/3}$ is a Hauptmodul for the level $3$ subgroup which is the kernel of the composite $\mathrm{PSL}_{2}(\mathbf{Z})\rightarrow\mathrm{PSL}_{2}(\mathbf{F}_{3})=A_{4}\rightarrow\mathbf{Z}/3\mathbf{Z}$. One final example is the function
(1.0.4) $h:=\frac{\lambda(\tau)(1-\lambda(\tau))}{16}=\left(\frac{\eta(\tau/2)\eta(2\tau)}{\eta(\tau)^{2}}\right)^{24}$
of level $\Gamma^{0}(2)\supset\Gamma(2)$; here the complete list of $n$ for which $h^{1/n}$ is either congruence modular or has bounded denominators are the divisors of $24$. The claim that $h^{1/n}$ has bounded denominators for $n|24$ is apparent from the product formula in Equation 1.0.4, and the claim that $h^{1/n}$ does not have bounded denominators for $n\nmid 24$ is an elementary exercise. We can directly compute when is $h^{1/n}$ congruence, as follows. By Kummer theory, the extension $\mathbf{C}(h^{1/n})$ of the function field $\mathbf{C}(h)$ of $\mathbf{H}/\Gamma^{0}(2)$ is Galois with Galois group $\mathbf{Z}/n\mathbf{Z}$. Since $h$ is nonvanishing on $\mathbf{H}$, this extension is unramified away from the cusps, and so gives rise to a homomorphism $\Gamma^{0}(2)\rightarrow\mathbf{Z}/n\mathbf{Z}$; the function $h^{1/n}$ is modular for the kernel $\Gamma$ of this homomorphism. Now $h^{1/n}$ is congruence if and only if the latter map factors through the congruence completion $\widehat{\Gamma^{0}(2)}$ of $\Gamma^{0}(2)$; here, $\Gamma^{0}(2)$ is considered as a subgroup of $\mathrm{PSL}_{2}(\mathbf{Z})$. In fact, one may compute that the abelianization of $\Gamma^{0}(2)$ is $\mathbf{Z}\oplus\mathbf{Z}/2\mathbf{Z}$ whereas the abelianization of $\widehat{\Gamma^{0}(2)}$ is $\mathbf{Z}/24\mathbf{Z}\oplus\mathbf{Z}/2\mathbf{Z}$. The other $\mathbf{Z}/2\mathbf{Z}$ extension corresponds to the congruence modular form $\sqrt{1-64h}=1-2\lambda$.
In a similar vein pertaining to the examples from the representation theory of vertex operator algebras, we prove in our closing § 7 the natural generalization of Theorem 1.0.1 to components of vector-valued modular forms for $\mathrm{SL}_{2}(\mathbf{Z})$, in particular resolving — in a sharper form, in fact — Mason’s unbounded denominators conjecture [Mas12, KM08] on generalized modular forms.
@@ 1.1.1. The Diophantine principle
Lemma 1.1.2 follows upon combining the following two observations, fixing some $1>\eta>R^{-1}$:
@@
We shall refer to the first inequality as a Liouville lower bound, following its use by Liouville in his proof of the lower bound $|\alpha-p/q|\gg 1/q^{n}$ for algebraic numbers $\alpha\neq p/q$ of degree $n\geq 1$. We shall refer to the second inequality as a Cauchy upper bound, following the example above where it comes from an application of the Cauchy integral formula. The first nontrivial generalization of Lemma 1.1.2 was Émile Borel’s theorem [Bor94]. Dwork famously used a $p$-adic generalization of Borel’s theorem in his $p$-adic analytic proof of the rationality of the zeta function of an algebraic variety over a finite field (see Dwork’s account in the book [DGS94, Chapter 2]). The simplest nontrivial statement of Borel’s theorem is that an integral formal power series $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ must already be a rational function as soon as it has a meromorphic representation as a quotient of two convergent complex-coefficients power series on some disc $D(0,R)$ of a radius $R>1$. The subject of arithmetic algebraization blossomed at the hands of many authors, including most prominently Carlson, Pólya, Robinson, Salem, Cantor, D. &\& G. Chudnovsky, Bertrandias, Zaharjuta, André, Bost, Chambert-Loir [CL02, BCL09, Ami75], [And04, § I.5], [And89, § VIII]. A simple milestone that we further develop in our § 2 is André’s algebraicity criterion [And04, Théorème 5.4.3], stating in a particular case that an integral formal power series $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ is algebraic as soon as the two formal functions $x$ and $f$ admit a simultaneous analytic uniformization — that means an analytic map $\varphi:(D(0,1),0)\to(\mathbf{C},0)$ such that the composition $f(\varphi(z))\in\mathbf{C}\llbracket z\rrbracket$ of holomorphic function germs also converges on the full disc $D(0,1)$, and such that $\varphi$ is sufficiently large in terms of conformal size, namely: $|\varphi^{\prime}(0)|>1$. For example, for any integer $m$, the algebraic power series $f=(1-m^{2}x)^{1/m}\in\mathbf{Z}\llbracket x\rrbracket$ admits the simultaneous analytic uniformization $x=\varphi(z)=(1-e^{Mz})m^{-2}$ and $f=f(\varphi(z))=e^{Mz/m}$, where the conformal size $|\varphi^{\prime}(0)|=M/m^{2}$ can clearly be made arbitrarily large by making a suitable choice of $M$.
@@ 1.1.3. Modularity and simultaneous uniformizations of $f$ and $\lambda$
Let us now explain the relevance of arithmetic holonomy rank bounds to the unbounded denominators conjecture. After reducing to weight $k=0$ as above, the functions $f=f(\tau)$ and $x:=\lambda(\tau)/16\in q+q^{2}\mathbf{Z}\llbracket q\rrbracket$ are algebraically dependent and share both (we assume) the property of integral Fourier coefficients at the cusp $q=0$. Let us assume for the purpose of this sketch that $f(\tau)\in\mathbf{Z}\llbracket q\rrbracket$ with $q=e^{\pi i\tau}$, i.e. that the cusp $i\infty$ has width dividing $2$. Then the formal inverse series expansion
$q=x+8x^{2}+84x^{3}+992x^{4}+\cdots\in x+x^{2}\mathbf{Z}\llbracket x\rrbracket$
of (1.0.2) has integer coefficients, expressing the identity $\mathbf{Z}\llbracket q\rrbracket=\mathbf{Z}\llbracket x\rrbracket$ of formal power series rings, and that formal substitution turns our integral Fourier coefficients hypothesis into an algebraic power series with integer coefficients: henceforth in this introductory sketch we switch to writing, by a mild and harmless notational abuse, simply $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ in place of $f(\tau)$ and $\lambda(q)$ in place of $\lambda(\tau)$. In the general case of arbitrary cusp width, which we need anyhow for the inner workings of our proof even if one is ultimately interested in the $\mathbf{Z}\llbracket q\rrbracket$ case, we will only have $f\in\mathbf{Z}[1/N]\llbracket x\rrbracket$ when we write out $f(\tau)$ as a power series in $x:=\sqrt[N]{\lambda/16}$ to accommodate the Puiseux series — but there is still a hidden integrality property which we can exploit. That leads to some mild technical nuance with the power series (2.0.2) — think of $t=q^{1/N}$, $x(t)=\sqrt[N]{\lambda(t^{N})/16}$ and $p(x)=x^{N}$ — in our refinement (2.0.3) of André’s theorem.
@@ 1.1.6. A finite local monodromy leads to an overconvergence
It turns out, and this is the key to our method and already answers André’s question in [And04, Appendix, A.5], that a different choice of $\varphi(z)$ allows one to arithmetically distinguish between these two cases (algebraic and transcendental), and to have the algebraicity of $f(x)$ recognized by André’s Diophantine criterion by way of an “overconvergence.” Suppose that $f(\tau)$ is a holomorphic modular function and $F(\tau)$ is a holomorphic modular form — concretely, let us take $F(\tau)$ to be the theta series of equation (1.1.5), — and let $f(x)$ and $F(x)$ respectively denote these functions as functions of $x=\lambda/16$, so $F(x)$ is given by equation (1.1.4). The common feature of these two functions $f(x)$ and $F(x)$ — coming respectively out of modular forms of weights 0 and $1$ — is that they both vary holonomically in $x\in\mathbf{C}\smallsetminus\{0,1/16\}$: they satisfy linear ODEs with coefficients in $\mathbf{Q}[x]$ and no singularities [^1] With nontrivial local monodromy. The precise definition is in 2.0.4. apart from the three punctures $x=0,1/16,\infty$ of $Y(2)=\mathbf{H}/\Gamma(2)$. The difference feature is that their respective local monodromies around $x=0$ are finite for the case of $f(x)$ (a quotient of $\mathbf{Z}/N$, with the order $N$ equal to the lowest common multiple of the cusp widths, or Wohlfahrt level [Woh64] of $f(x)$); and infinite for the case of $F(x)$ (isomorphic to $\mathbf{Z}$, corresponding more particularly to the fact that this particular hypergeometric function acquires a $\log{x}$ term after an analytic continuation around a small circle enclosing $x=1/16$). If now we perform the variable change $x\mapsto x^{N}$, redefaulting to $x:=\sqrt[N]{\lambda(q^{N})/16}$, that resolves the $N$-th root ambiguity in the formal Puiseux branches of $f(x)$ at $x=0$, and the resulting algebraic power series $f(x^{N})\in\mathbf{Z}\llbracket x^{N}\rrbracket\subset\mathbf{Z}\llbracket x\rrbracket$ has turned holonomic on $\mathbf{P}^{1}\smallsetminus\{16^{-1/N}\mu_{N},\infty\}$: singularities only at $16^{-1/N}\mu_{N}\cup\{\infty\}$ (but not at $x=0$: this key step of exploiting arithmetic algebraization is the same as in Ihara’s arithmetic connectedness theorem [Iha94, Theorem 1], which together with Bost’s extension [Bos99] to arithmetic Lefschetz theorems have in equal measure been inspirational for our whole approach to the unbounded denominators conjecture). Since $\lambda:D(0,1)\to\mathbf{C}\smallsetminus\{1\}$ has fiber $\lambda^{-1}(0)=\{0\}$, the function $\varphi(z):=\sqrt[N]{\lambda(z^{N})/16}:D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ is still holomorphic on the unit disc $|z|<1$, and under this tautological choice, both functions $f(x^{N})$ and $F(x^{N})$ continue to be at the borderline of André’s algebraicity criterion: $|\varphi^{\prime}(0)|=1$.
But if instead of the tautological simultaneous uniformization we take
$\varphi:D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$
@@
to be the universal covering map (pointed at $\varphi(0)=0$), then either by a direct computation with monodromy, or by Cauchy’s analyticity theorem on the solutions of linear ODEs with analytic coefficients and no singularities in a disc, we have both function germs $x:=\varphi(z)$ and $f(x):=f(\varphi(z))$ holomorphic, hence convergent, on the full unit disc $D(0,1)$. In contrast, now $F(\varphi(z))$ converges only up to the “first” nonzero fiber point in $\varphi^{-1}(0)\smallsetminus\{0\}$, giving a certain radius rather smaller than $1$. We must have the strict lower bound $|\varphi^{\prime}(0)|>1$, because the preceding unit-radius holomorphic map $\sqrt[N]{\lambda(z^{N})/16}:D(0,1)\to\mathbf{C}\smallsetminus 16^{-1/N}\mu_{N}$ has to factorize properly through the universal covering map. Indeed in Theorem 5.1.4, using an explicit description by hypergeometric functions of the multivalued inverse of the universal covering map of $\mathbf{C}\smallsetminus\mu_{N}$ based on Poincaré’s ODE approach [Hem88] to the uniformization of Riemann surfaces, we find an exact formula for this uniformization radius in terms of the Euler Gamma function. [^2] André pointed out to us that this explicit formula has previously been obtained by Kraus and Roth, see [KR16, Remark 5.1]. See also [Gol69, § III.1]. Hence the algebraicity of $f(x)$ gets witnessed by André’s criterion; and the formal new result that we get already at this opening stage (see Theorem 7.2.1) is that any integral formal power series solution $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ to a linear ODE $L(f)=0$ without singularities on $\mathbf{P}^{1}\smallsetminus\{0,1/16,\infty\}$ is in fact algebraic as soon as the linear differential operator $L$ has a finite local monodromy $\mathbf{Z}/N$ around the singular point $x=0$. More than this: the quantitative Corollary 2.0.5 proves that the totality of such $f(x)\in\mathbf{Z}\llbracket x\rrbracket$ at a given $N$ span a finite-dimensional $\mathbf{Q}(x)$-vector space, and gives an upper bound on its dimension as a function of the Wohlfahrt level parameter $N$. Now since a (noncongruence) counterexample $f(\tau)\in\mathbf{Z}\llbracket q\rrbracket$ to Theorem 1.0.1 would not exist on its own but spawn a whole sequence $f(p\tau)\in\mathbf{Z}\llbracket q\rrbracket$ of $\mathbf{Q}(x)$-linearly independent counterexamples at growing Wohlfahrt level $N\mapsto Np$, our idea is to measure up the supply of these putative (fictional) counterexamples alongside the congruence supply at a gradually increasing level until together they break the quantitative bound (2.0.3) supplied by our arithmetic holonomy Theorem 2.0.1.